# Class-Dependent Perturbation Effects in Evaluating Time Series Attributions

### v1

@@ 5.4 Evaluating Constant Perturbation Values
Table 5 shows the class-adjusted DS for the constant perturbation strategies, revealing some consistent patterns in the datasets while highlighting important domain-specific variations.
Moderate positive perturbation values between 0 and 1 often achieve the best performance, although effectiveness varies by dataset and classifier architecture. FordA and FordB demonstrate the best performance with 0.5 perturbation. Wafer is inconsistent, with different perturbation constants working best for different attribution methods. ElecDev exhibits the best performance with lower perturbation values between -0.5 and 0.5, though its multiclass nature introduces additional complexity in interpreting class-adjusted metrics.
Interestingly, negative perturbation values consistently produce negative class-adjusted DS scores for most datasets, indicating asymmetric class responses to the perturbation direction. However, this effect varies by attribution method, classifier, and dataset. For example, SG and GS on ElecDev only show positive adjusted DS scores. These results suggest that the effectiveness of perturbation-based evaluation depends on the alignment between perturbation values and the learned class representations.
@@ 5.5 Synthesis and Implications
These findings have important implications for XAI evaluation methodology. The observed class-dependent effects indicate that aggregate performance metrics benefit from additional analysis that considers class-specific behaviors. When designing evaluation protocols, researchers may want to examine whether their chosen perturbation strategies show consistent effectiveness across classes, particularly in contexts with class imbalance or when working with new domains and model architectures.
Table 5: Class-adjusted mean degradation scores ($\text{DS}_{c}$ with $\alpha=1$) for constant perturbation strategies across datasets, models, and attribution methods. $\text{DS}_{c}$ measures both attribution quality and cross-class consistency of perturbations. Positive values indicate good attribution quality with consistent behavior across classes, while lower values suggest either poor attribution quality or inconsistent performance between classes. Bold and italic values highlight the highest and second-highest values per dataset and model-attribution combination.
ResNet InceptionTime
Dataset Perturbation GR IG SG GS FO GR IG SG GS FO
FordA -2 -0.05 -0.07 -0.01 -0.07 -0.08 -0.02 -0.08 -0.00 -0.04 -0.08
-1.5 -0.07 -0.09 -0.01 -0.09 -0.09 -0.03 -0.10 -0.00 -0.06 -0.11
-1 -0.10 -0.10 -0.01 -0.10 -0.09 -0.05 -0.09 0.00 -0.05 -0.09
-0.5 -0.13 -0.01 -0.00 -0.02 0.00 -0.14 0.00 0.00 -0.00 -0.00
0 -0.08 0.00 0.00 0.00 0.01 -0.20 0.00 0.00 -0.00 -0.00
0.5 -0.03 0.00 0.00 0.00 0.01 -0.04 0.00 0.00 0.00 0.00
1 -0.01 0.00 0.00 0.00 0.00 -0.01 0.00 -0.00 0.00 0.00
1.5 -0.01 0.00 0.00 0.00 0.00 -0.00 0.00 -0.00 0.00 0.00
2 -0.01 0.00 0.00 0.00 0.00 -0.00 0.00 -0.00 0.00 0.00
FordB -2 -0.01 -0.04 0.00 -0.04 -0.02 -0.00 -0.05 -0.00 -0.04 -0.06
-1.5 -0.02 -0.05 0.00 -0.05 -0.02 -0.01 -0.06 -0.00 -0.05 -0.07
-1 -0.03 -0.03 0.00 -0.04 -0.01 -0.02 -0.06 -0.02 -0.06 -0.08
-0.5 -0.04 0.00 0.00 -0.00 0.00 -0.05 -0.02 -0.06 -0.02 -0.02
0 -0.04 0.01 0.00 0.01 0.01 -0.16 0.01 -0.13 0.01 0.02
0.5 -0.02 0.01 -0.00 0.01 0.02 -0.10 0.02 -0.02 0.02 0.04
1 -0.00 0.00 -0.00 0.00 0.01 -0.07 0.02 0.01 0.01 0.03
1.5 -0.00 0.00 -0.00 0.00 0.00 -0.05 0.01 0.01 0.01 0.01
2 -0.00 0.00 -0.00 0.00 0.00 -0.03 0.00 0.00 0.00 0.01
Wafer -2 -0.03 -0.05 0.00 -0.05 -0.13 0.00 -0.12 -0.00 -0.15 -0.37
-1.5 -0.09 -0.17 0.00 -0.19 -0.29 0.00 -0.17 -0.00 -0.25 -0.42
-1 -0.20 -0.20 0.00 -0.27 -0.31 0.01 -0.23 -0.01 -0.31 -0.34
-0.5 -0.14 -0.05 -0.00 -0.08 -0.05 0.02 -0.18 -0.01 -0.22 0.02
0 -0.05 0.01 -0.00 0.01 0.01 0.01 0.01 -0.02 0.04 0.02
0.5 0.01 0.01 0.00 0.01 0.02 0.01 0.01 -0.03 0.05 0.05
1 0.02 0.00 0.00 0.00 0.02 -0.01 0.00 -0.03 0.03 0.06
1.5 0.03 -0.00 -0.00 -0.00 0.01 0.00 0.00 -0.02 0.02 0.04
2 0.02 -0.01 -0.00 -0.00 -0.00 -0.00 0.00 -0.01 0.01 0.03
ElecDev -2 -0.05 -0.04 -0.03 -0.04 -0.05 0.00 -0.03 0.04 0.01 -0.11
-1.5 -0.02 -0.04 -0.01 -0.03 -0.03 0.02 -0.01 0.05 0.02 -0.07
-1 0.02 -0.07 0.00 -0.03 -0.01 0.02 0.00 0.05 0.04 -0.02
-0.5 0.00 -0.03 0.02 0.01 0.05 0.05 0.06 0.08 0.09 0.10
0 -0.01 0.09 0.03 0.06 0.23 0.05 0.12 0.04 0.10 0.16
0.5 0.00 0.15 0.10 0.13 0.27 0.01 0.11 0.02 0.11 0.14
1 0.01 0.11 0.08 0.10 0.19 0.01 0.11 0.03 0.09 0.12
1.5 0.01 0.09 0.08 0.09 0.14 0.02 0.10 0.03 0.07 0.06
2 0.00 0.07 0.08 0.07 0.08 0.02 0.04 0.04 0.03 -0.00
@@ 6 Conclusion and Future Work
This paper presents a systematic investigation of class-dependent effects in perturbation-based evaluation of feature attributions for time series classification. Through empirical evaluation across four datasets, five attribution methods, and multiple perturbation strategies, we demonstrate that perturbation-based evaluation methods can exhibit class-specific behaviors that warrant careful consideration in validation procedures.
Our investigation yields three main findings. First, we show that perturbation strategies can exhibit notable class-dependent effects, where attribution validation may show varying effectiveness across different classes. Second, we demonstrate that perturbation strategies that show strong aggregate performance may derive their effectiveness partially from class-specific effects rather than solely from attribution quality. When controlling for class-dependent behaviors through our proposed penalty framework, we see that most of their effects come from strong performance for only one class. Third, our analysis reveals that the best perturbation values, including alternatives to the commonly used zero-value substitution, vary substantially across datasets and model architectures. This indicates that perturbation effectiveness can depend heavily on learned model representations rather than intrinsic attribution quality.
These findings have important implications for XAI evaluation methodology. The presence of class-dependent effects suggests that aggregate performance metrics benefit from additional analysis, particularly in imbalanced datasets where dominant classes might influence overall performance metrics. Furthermore, the variation in optimal perturbation values between different experimental configurations indicates that evaluation protocols may benefit from dataset-specific calibration. To address these challenges, we introduce a class-adjusted penalty framework applicable to any aggregate XAI evaluation metric, enabling researchers to investigate potential class-specific patterns without additional computational overhead. Beyond aggregate analysis, we recommend examining individual instances with multiple perturbation strategies designed to push model predictions toward different classes. If an attribution’s effect can be validated using at least one such strategy, this can suggest that the attribution itself is sound, even if other strategies prove ineffective due to class-specific behaviors.
Several limitations of our study should be noted. Our investigation focuses on only four datasets and two classifier architectures, and while we observe consistent patterns across multiple perturbation strategies, these findings may not generalize to all contexts. For instance, the observed patterns in the responses of the classifiers to perturbations may reflect specific characteristics of our experimental setup, such as the way the classifiers were trained, rather than general properties of the evaluation approach. However, our results align well with previous studies in the literature, suggesting broader applicability of our methodology. We encourage researchers to validate and extend these findings using their own datasets and experimental configurations.
@@
Future research directions emerge from these findings. One promising avenue involves developing perturbation strategies that systematically account for class-specific model behaviors, particularly focusing on methods that can push predictions toward different classes. This approach could lead to more comprehensive evaluation metrics that consider the directionality of perturbation effects. Additionally, investigating methods to adaptively select perturbation strategies for individual instances could improve evaluation effectiveness, as our results suggest that different instances may require different perturbation approaches to effectively validate their attributions. Understanding this interaction between instance characteristics and perturbation effectiveness could provide insight into designing more robust evaluation frameworks.

### v2

@@ 5.4 Evaluating Constant Perturbation Values
Table 5 shows the class-adjusted DS for the constant perturbation strategies, revealing significant domain-specific variations. The effectiveness of constant perturbations varies significantly by dataset and classifier architecture. FordA and FordB often show the best performance with 0.5 perturbation. Wafer demonstrates inconsistent responses, where optimal perturbation constants vary across attribution methods. ElecDev exhibits the best performance with a perturbation constant of 0.5 for ResNet, and different constants between -0.5 and 0.5 for InceptionTime, though its multiclass nature introduces additional complexity in interpreting class-adjusted metrics.
Interestingly, negative perturbation values generally yield negative class-adjusted scores for most datasets, indicating asymmetric class responses to the perturbation direction. However, this effect varies by attribution method, classifier, and dataset. For example, GR, SG and GS on ElecDev exclusively show positive scores for negative perturbation constants. These results suggest that the effectiveness of perturbation-based evaluation depends on the alignment between perturbation values and the learned class representations in the feature space.
Table 5: Class-adjusted mean degradation scores ($\text{DS}_{c}$ with $\alpha=1$) for constant-value perturbation strategies across datasets, models, and attribution methods. $\text{DS}_{c}$ measures both attribution quality and cross-class consistency of perturbations. Positive values indicate good attribution quality with consistent behavior across classes, while lower values suggest either poor attribution quality or inconsistent performance between classes. The largest values for each dataset and model-attribution column pair are highlighted in bold.
ResNet InceptionTime
Dataset Perturbation GR IG SG GS FO GR IG SG GS FO
FordA -2 -.045 -.068 -.008 -.067 -.083 -.016 -.077 .000 -.042 -.083
-1.5 -.066 -.093 -.012 -.092 -.094 -.027 -.100 .000 -.056 -.106
-1 -.098 -.102 -.013 -.104 -.086 -.054 -.091 .000 -.054 -.091
-0.5 -.129 -.011 .000 -.016 .001 -.136 .000 .001 .000 .000
0 -.082 .003 .001 .002 .005 -.197 .000 .001 .000 -.001
0.5 -.032 .005 .001 .004 .010 -.040 .001 .000 .001 .002
1 -.012 .003 .001 .003 .004 -.009 .001 .000 .000 .002
1.5 -.009 .001 .001 .001 .001 -.005 .001 .000 .000 .001
2 -.009 .000 .000 .000 .001 -.004 .000 .000 .000 .000
FordB -2 -.011 -.038 .001 -.043 -.015 -.004 -.045 .000 -.039 -.056
-1.5 -.017 -.045 .001 -.054 -.018 -.007 -.060 -.002 -.054 -.074
-1 -.026 -.033 .001 -.043 -.009 -.017 -.063 -.017 -.060 -.076
-0.5 -.040 .000 .001 .000 .003 -.054 -.016 -.062 -.023 -.020
0 -.044 .006 .000 .005 .010 -.159 .013 -.134 .015 .015
0.5 -.024 .008 .000 .007 .018 -.096 .022 -.021 .020 .039
1 -.005 .003 -.003 .003 .006 -.069 .015 .009 .013 .028
1.5 -.001 .001 -.002 .001 .002 -.053 .009 .006 .008 .013
2 -.001 .000 -.001 .000 .000 -.034 .005 .003 .004 .005
Wafer -2 -.035 -.054 .001 -.053 -.125 .001 -.116 -.002 -.155 -.374
-1.5 -.087 -.166 .001 -.191 -.289 .002 -.169 -.004 -.250 -.423
-1 -.204 -.202 .000 -.267 -.309 .006 -.226 -.006 -.314 -.337
-0.5 -.142 -.048 -.004 -.075 -.048 .018 -.177 -.006 -.224 .018
0 -.047 .007 -.001 .011 .013 .012 .009 -.023 .036 .016
0.5 .008 .008 .000 .008 .017 .005 .014 -.029 .047 .047
1 .025 .002 .001 .003 .016 -.014 .004 -.027 .031 .056
1.5 .031 -.003 -.001 -.001 .006 .000 .003 -.016 .017 .040
2 .019 -.006 -.003 -.003 -.001 -.001 .003 -.013 .014 .030
ElecDev -2 -.055 -.040 -.032 -.037 -.054 .004 -.028 .044 .007 -.114
-1.5 -.023 -.039 -.010 -.031 -.029 .016 -.010 .048 .024 -.074
-1 .017 -.069 .000 -.027 -.005 .021 .003 .054 .036 -.021
-0.5 .002 -.028 .018 .011 .049 .054 .055 .075 .092 .104
0 -.010 .089 .031 .063 .226 .053 .118 .038 .104 .162
0.5 .005 .147 .100 .129 .269 .013 .106 .019 .106 .144
1 .006 .114 .076 .098 .193 .005 .112 .027 .094 .123
1.5 .005 .087 .078 .086 .139 .015 .098 .032 .066 .063
2 .003 .067 .078 .068 .078 .024 .041 .036 .029 -.005
@@ 6 Conclusion and Future Work
This paper presents a systematic investigation of class-dependent effects in perturbation-based evaluation of feature attributions for time series classification. Through empirical evaluation across four datasets, five attribution methods, and multiple perturbation strategies, we demonstrate that perturbation-based evaluation methods can exhibit class-specific behaviors that warrant careful consideration in validation procedures. Namely, we show that: (1) perturbation effectiveness can vary substantially between classes; (2) strategies with stronger aggregate performance often exhibit more pronounced class-dependent effects; and (3) optimal perturbation strategies can vary considerably across datasets and model architectures, suggesting that perturbation effectiveness is influenced by specific characteristics of learned model representations and that domain-specific calibration may be necessary.
We recommend several evaluation approaches to address potential class-dependent effects. First, supplementing aggregate metrics with more detailed class-stratified analysis can help identify when dominant classes disproportionately influence results. Second, applying penalty frameworks like the one proposed in this study quantifies class-dependent effects without additional computational burden, allowing researchers to systematically compare attribution methods while accounting for class biases. Third, evaluating individual instances with multiple perturbation strategies targeting different class predictions may help validate attribution robustness beyond the limitations of any single approach.
The generalizability of our findings faces two primary constraints. Our investigation encompasses only four datasets and two classifier architectures, potentially limiting broader applicability. Additionally, the observed perturbation response patterns may reflect specific characteristics of our experimental configuration rather than fundamental properties of the evaluation approach. Nevertheless, the alignment with previous literature suggests wider relevance of our methodology.
@@
Future research directions emerge from these findings. One promising avenue involves developing perturbation strategies that systematically account for class-specific model behaviors, particularly focusing on methods that can push predictions toward different classes. Additionally, investigating methods to adaptively select perturbation strategies for individual instances could improve evaluation effectiveness, as our results suggest that different instances may require different perturbation approaches to effectively validate their attributions.
Appendix 0.A Dataset Visualizations and Class Distributions
Table 6: Class distribution across datasets. Values represent the proportion of samples in each class for training and test sets.
Train Test
Dataset C0 C1 C2 C3 C4 C5 C6 C0 C1 C2 C3 C4 C5 C6
FordA .513 .487 .516 .484
FordB .512 .488 .495 .505
Wafer .097 .903 .108 .892
ElecDev .081 .250 .095 .165 .270 .057 .082 .087 .254 .098 .151 .242 .096 .072
(a) FordA dataset
(b) FordB dataset
(c) Wafer dataset
Figure 6: Representative samples from the three binary classification datasets used in this study. For each dataset, three randomly selected instances per class are shown to illustrate the characteristic patterns and variability between classes. If datasets had labels not starting from 0, they were remapped to start from 0 (ranging from 0 to $C-1$), preserving the existing order.
Figure 7: Representative samples from the one multiclass classification dataset used in this study, ElecDev. Three randomly selected instances per class are shown to illustrate the characteristic patterns and variability between classes. If datasets had labels not starting from 0, they were remapped to start from 0 (ranging from 0 to $C-1$), preserving the existing order.